OpenStudy (anonymous):

Suppose the population of a country in 1985 was 145 million. In 1995 it was 190 million. Use the formula P = Ae kt to setup an expression to estimate the population of the country in 2005

OpenStudy (anonymous):

Let t be the year in question, then P(t) = Ae^k(t-1985) P(1985) = Ae^0=> A = 145 P(1995)=145e^k(1995-1985)=> 190=145e^10k=> 10k=ln(190/145)=> k = 0.027 P(2005) = 145e^0.027(2005-1985)=> = 145e^0.027(20)=> = 249 million approximately.